Two 2.0 x 2.0 cm square metal plates, spaced 0.50 mm apart are connected to a 100 V battery.

What is the Electric field strength in the capacitor?
Answer #1

Electric field, E = V/d
Where V is the voltage in the battery and d is the separation between the metal plates.
V = 100 V
d = 0.5 mm
= 0.5 * 10-3 m

E = 100 / (0.5 * 10-3)
= 2.0 * 105 V/m